# Network Oblivious Transfer via Noisy Channels: Limits and Capacities

### v2

@@ III-B2 OT capacity
As we stated before, a few research works consider the problem of OT from the information theory perspective. Nascimento and Winter [6, 13] simplified the problem of a general noisy correlation (a point-to-point channel) by reducing it to a Slightly Unfair Noisy Symmetric Basic Correlation (SU-SBC). They demonstrated that any non-perfect noisy point-to-point channel or correlation can be transformed into a Slightly Unfair Noisy Channel/Correlation (SUNC/SUCO), and that any SUNC/SUCO can be used to implement a specific SU-SBC. Ultimately, they showed that in this reduced framework, it is possible to achieve $\binom{m}{1}\text{-OT}^{k}$ (11-out-of-$m$ OT with strings’ length equal $k$) at a positive rate, assuming that the sender behaves in an honest-but-curious manner. These papers are significant for several reasons:
@@
Ahlswede and Csiszár analyzed the problem under the honest-but-curious model [5]. They derived a general upper bound for the OT capacity of a point-to-point DMC, which aligns with the lower bound established by Nascimento and Winter [6]. Furthermore, when reduced to a specific erasure channel, they demonstrated a lower bound on the OT capacity. Thus, it remains uncertain what the exact OT capacity of a noisy DMC is in general. In practice, for general channels, a potential way is to first convert the channel into a Generalized Erasure Channel (GEC) via alphabet extension and erasure emulation, followed by the application of a general construction for GEC. As a starting point, we begin with a protocol by Ahlswede and Csiszár originated from the general two-party protocol of Section II:
@@
Two-Party OT Protocol [5]: Consider the following two-party secure computation in the sense of OT: Alice has two strings $M_{0}$ and $M_{1}$ and aims to send them over the noisy point-to-point channel $\mathcal{W}:\mathcal{X}\rightarrow\mathcal{Y}$ to Bob. Bob has to choose one of them by inputting a bit of $Z\in\{0,1\}$ to the channel. Alice should be unaware of the unselected string, while Bob has only one string at the end of the protocol $(M_{Z})$. Suppose the main channel is an erasure channel assisted by a noiseless channel with unlimited capacity. The OT capacity in this setup is given by $\min(p,1-p)$, where $p$ is the erasure probability [5, 14]. Let $r<\min(p,1-p)$. The protocol by Ahlswede and Csiszár [5], based on a technique originally introduced for a BSC in [9, Sec. 6.4], proceeds as follows: Alice starts by transmitting a sequence $\mathbf{X}=X^{n}\sim\text{Bernoulli}(\frac{1}{2})$ of i.i.d. bits over the channel. Bob observes the channel output $\mathbf{Y}$. Let $E$ denote the set of indices at which $\mathbf{Y}$ is erased, and let $\overline{E}$ represent the set of indices where $\mathbf{Y}$ is not erased. If $|E|<nr$ or $|\overline{E}|<nr$, Bob aborts the protocol, as there are not enough erased or unerased bits to complete the protocol. From $\overline{E}$, Bob randomly selects a subset $S_{Z}$ of cardinality $nr$. From $E$, Bob randomly picks a subset $S_{\overline{Z}}$, also of size $nr$. Bob then shares the sets $S_{0}$ and $S_{1}$ with Alice via the public channel, where $S_{0}$ and $S_{1}$ are either $S_{Z}$ and $S_{\overline{Z}}$, respectively or vice versa. Alice cannot determine which of $S_{0}$ and $S_{1}$ corresponds to $E$ (erased positions) and which to $\overline{E}$ (non-erased positions) due to the independent nature of channel erasures. Using $S_{0}$ and $S_{1}$, Alice computes keys $\mathbf{X}|_{S_{0}}$ and $\mathbf{X}|_{S_{1}}$ and employs these keys to encrypt her strings, which she sends to Bob over the public channel as $M_{0}\oplus\mathbf{X}|_{S_{0}}$ and $M_{1}\oplus\mathbf{X}|_{S_{1}}$. Bob, who only knows the sequence corresponding to $\overline{E}$, can decrypt only one of these encrypted strings, depending on whether $S_{Z}=S_{0}$ or $S_{Z}=S_{1}$. This enables Bob to retrieve one of the two keys, $M_{Z}$, while he learns nothing about the other key, $M_{\overline{Z}}$. If $X$ is not uniformly distributed over $\{0,1\}$, the strings $\mathbf{X}|_{S_{j}},j\in\{0,1\}$ are not directly suitable as encryption keys. They need to be transformed into binary strings of length $k<nr$ with a distribution approximately uniform over $\{0,1\}^{k}$. It is well-known that for any $\delta>0$, when $n$ is large, there exists a mapping $\kappa:\{0,1\}^{nr}\rightarrow\{0,1\}^{k}$ with $k=n(H(X)-\delta)$ such that $k-H(\kappa(X^{n}))$ is exponentially small.
@@
In [32], the author extends the above protocol to pairwise oblivious transfer over a noiseless binary adder channel involving two senders and one receiver, assuming they are non-colluding and honest-but-curious. Each sender has two strings, and Bob has to choose one string from each sender while the unselected strings are hidden from his view. In this system, the output is defined as the sum of the inputs, $Y=X_{1}+X_{2}$, commonly referred to as the Binary Erasure Multiple Access Channel (BE-MAC). This channel uniquely determines the inputs except when they differ, in which case one can not identify the inputs with certainty, effectively resulting in an erasure. Specifically, erasures occur in two out of four possible input scenarios. The OT capacity of this channel is shown to be $R_{1}+R_{2}\leq\max_{P_{X_{1}}P_{X_{2}}}H(X_{1},X_{2}|Y)=\frac{1}{2}$.
@@ IV System Model
A key concept when analyzing noisy channels is the idea of redundant symbols [33]. We have the following definition for DM-MACs, presented in [6] for the point-to-point channel.
@@
As proved in [6] for point-to-point channels, a perfect DM-MAC (even after removing the redundancy) cannot be used for oblivious transfer, even against passive adversaries. This relates to the concept of noise and the emergence of noisy resources for cryptographic intents. We know that the noise produces uncertainty or entropy. It is the unpredictable aspect of a system that adds complexity. Conceptually, it can be seen as the manifestation of disorder or randomness in communication systems, emphasizing the non-perfect nature of the real world. Noise plays a crucial role in securing communication by hindering an eavesdropper’s ability to extract meaningful information from the transmitted message. In a noisy channel, the inherent noise limits the amount of information an unauthorized party can access, regardless of their computational power. This concept is central to Wyner’s wiretap channel model [34], which demonstrates how noise can be leveraged to ensure that the legitimate receiver decodes the message accurately. In contrast, experiencing additional noise, the eavesdropper cannot gather sufficient information to reconstruct the message. As is clear, a perfect channel can be simulated by a noiseless channel where the input(s) can be obtained with certainty from the channel output(s). Obviously, such channels cannot be used for cryptographic intents with unconditional security (information-theoretic secrecy).
As is proved in [14, 5], the OT capacity of the point-to-point erasure channel (BEC) with erasure probability $\frac{1}{2}$ is equal $\frac{1}{2}$. Here, we want to investigate whether the OT is possible over a special BE-MAC. We introduce the channel as a correlation between the senders (Alice-1, Alice-2) and the receiver (Bob). Before that, we delve deeper into the unfairness in the channel/correlation model. Damgård et. al. [18], introduced unfairness so that an unfair player could change the communication channel parameters within a certain range. In [6], this concept is limited so that an unfair player who deviates from the channel statistics in $\delta n$ positions will be caught by the other party with probability $\geq 1-C_{1}\exp(-C_{2}\delta^{2}n)$, where $C_{1}$ and $C_{2}$ are two small positive numbers. There are two senders in our channel model. The concept of unfairness can also be extended to the MAC model. To control the fairness of other players in the last $n$ rounds, all players have access to a test unit. When both senders send a pair symbol to Bob over the channel, Bob will ask for the input symbols with probability $\frac{1}{2}$. He will tell both senders his output when he has received the senders’ response. If Alice-$i$ is unfair, she tells her input wrong; if Bob is unfair, he tells his output wrong. The test between players is to check out the samples after $n$ uses of the channel for joint typicality relative to $P_{X_{1}X_{2}Y}$.
@@
Our protocol to achieve $\text{SU-SBC}_{p,W}$ closely follows the approach outlined in [6, 8]. It uses the source coding protocol $P_{X_{1}X_{2}Y}$ twice for two independent realizations. Let:
@@
Bob cannot behave unfairly beyond what is provided by the SUCO $P_{X_{1}X_{2}Y}$. However, the senders can introduce bias by attempting actions such as repeating unfair pairs $(x_{1},x_{2})(x_{1},x^{\prime}_{2}),(x_{1},x_{2})(x^{\prime}_{1},x_{2})$ or $(x_{1},x_{2})(x^{\prime}_{1},x^{\prime}_{2})$ for a good pair $i\in\mathcal{L}$. Suppose the senders persist with such strategies for $\delta n$ times (where $\delta>0$). In that case, Bob can detect their bias by approximating a typicality test, using Lemma 6 and an extended version of [6, Lemma 6]. This aligns with the definition of $\text{SBC}_{p,W}$.
@@ V-A1 Information-Theoretic Formulation of OT
In the foundational works [35, 36], the information-theoretic security definition for the receiver’s privacy requires that the sender’s final view, denoted by $U=(U^{\prime},R_{A},\mathbf{C})$, be statistically independent of the receiver’s choice bit $z\in\{0,1\}$. Here, $U^{\prime}$ represents the sender’s initial state (including inputs), $R_{A}$ denotes the sender’s private randomness, and $\mathbf{C}$ captures the complete public communication between the sender and receiver. However, as highlighted by Crépeau in [37], such a strong condition is generally unachievable in many practical settings. In particular, when there is a known dependency between the parties’ inputs, the sender’s view will unavoidably correlate with the receiver’s input. This is because the sender’s input $X$, which is inherently part of her view, may be statistically dependent on the receiver’s input $Z$. To address this, a more appropriate formulation requires that the sender’s view be conditionally independent of the receiver’s input, given the sender’s input. Formally, this is expressed via the Markov chain $(U^{\prime},R_{A},\mathbf{C})\rightarrow X\rightarrow Z$, which implies the conditional mutual information satisfies $I(U^{\prime},R_{A},\mathbf{C};Z\mid X)=0$. Nevertheless, Ahlswede and Csiszár [5] adopt a slightly relaxed criterion, aligning with the interpretation in [35, 36], where the privacy condition is written as $I(M_{0},M_{1},X^{n},R_{A},\mathbf{C};Z)\rightarrow 0$. This formulation is weaker, as it allows for some residual dependence under the assumption that it diminishes asymptotically.
@@
Formulating the security definition for the sender presents additional challenges beyond those encountered in defining receiver security [35, 36]. Several issues commonly arise in this context:
@@
Notably, in [39, 38], the model constrains a malicious receiver to alter their input in a purely deterministic manner. That is, the effective input $Z^{\prime}$—representing the bit ultimately learned by the receiver—is required to be a deterministic function of their original input $Z$. This stands in contrast to the ideal functionality of oblivious transfer, where a malicious receiver is permitted to choose $Z^{\prime}$ probabilistically, potentially based on arbitrary strategies. Such a deterministic restriction limits the adversary’s capabilities in the simulation, potentially weakening the security guarantee for the sender in realistic adversarial settings.
@@
In scenarios where a malicious receiver may alter their effective input from $z$ to $z^{\prime}$ (with $z,z^{\prime}\in\{0,1\}$), [40] allows $Z^{\prime}$ to depend on the sender’s input $X$. Such dependency is not permitted in the ideal model, where parties’ inputs and outputs are generated independently of each other through a trusted functionality. Furthermore, the security definition requires that the view of the malicious receiver—denoted by $(V^{\prime},R_{B},Y^{n},\mathbf{C})$—be conditionally independent of the sender’s input $X$, given the receiver’s original input $Z$ and their output $M_{Z^{\prime}}$. This corresponds to the Markov chain: $X\rightarrow(Z,M_{Z^{\prime}})\rightarrow(V^{\prime},R_{B},Y^{n},\mathbf{C}),$ which implies: $I(V^{\prime},R_{B},Y^{n},\mathbf{C};X\mid Z,M_{Z^{\prime}})=0$.
@@
However, a more accurate formulation would account for the potential dependence of the receiver’s strategy on $Z^{\prime}$, leading to the stronger Markov condition: $X\rightarrow(Z,Z^{\prime},M_{Z^{\prime}})\rightarrow(V^{\prime},R_{B},Y^{n},\mathbf{C})$. Ahlswede and Csiszár [5] propose an alternative criterion under the honest-but-curious model—where Alice and Bob follow the protocol faithfully—that requires:
@@
In the context of OT, a malicious receiver may alter their effective input from $z$ to $z^{\prime}$ for $z,z^{\prime}\in\{0,1\}$. In [40], the variable $Z^{\prime}$—representing the receiver’s effective choice—may depend on the honest sender’s input $X$, which deviates from the ideal functionality where such a dependency is disallowed. Moreover, the receiver’s view, denoted by $(V^{\prime},R_{B},Y^{n},\mathbf{C})$, is required to be conditionally independent of the sender’s input $X$, given the receiver’s original input $Z$ and the corresponding output $M_{Z^{\prime}}$. This condition corresponds to the Markov chain: $X\rightarrow(Z,M_{Z^{\prime}})\rightarrow(V^{\prime},R_{B},Y^{n},\mathbf{C})$, or equivalently, $I(V^{\prime},R_{B},Y^{n},\mathbf{C};X\mid Z,M_{Z^{\prime}})=0$. However, a more accurate and robust formulation accounts for the dependency of the adversary’s strategy on the full pair $(Z,Z^{\prime})$, implying the stronger Markov chain: $X\rightarrow(Z,Z^{\prime},M_{Z^{\prime}})\rightarrow(V^{\prime},R_{B},Y^{n},\mathbf{C})$. Ahlswede and Csiszár [5] propose a different formulation in the honest-but-curious setting, where both parties follow the protocol faithfully. In this model, Alice’s input is the message pair $(M_{0},M_{1})$, and Bob’s input is $Z$. Their security criterion requires that: $I(Z,R_{B},Y^{n},\mathbf{C};X)\triangleq I(Z,R_{B},Y^{n},\mathbf{C};M_{\overline{Z}})$, and they study its asymptotic behavior under the i.i.d. assumption. While this approach simplifies the analysis for semi-honest participants, it is too permissive under malicious adversaries and may inadvertently allow protocols with security vulnerabilities.

### v4

@@ III-B2 OT capacity
As we stated before, a few research works consider the problem of OT from the information theory perspective. Nascimento and Winter [Winter1, Winter3] simplified the problem of a general noisy correlation (a point-to-point channel) by reducing it to a Slightly Unfair Noisy Symmetric Basic Correlation (SU-SBC). They demonstrated that any non-perfect noisy point-to-point channel or correlation can be transformed into a Slightly Unfair Noisy Channel/Correlation (SUNC/SUCO), and that any SUNC/SUCO can be used to implement a specific SU-SBC. Ultimately, they showed that in this reduced framework, it is possible to achieve $\binom{m}{1}\text{-OT}^{k}$ (11-out-of-$m$ OT with strings’ length equal $k$) at a positive rate, assuming that the sender behaves in an honest-but-curious manner. These papers are significant for several reasons:
@@
Ahlswede and Csiszár analyzed the problem under the honest-but-curious model [Rudolf1]. They derived a general upper bound for the OT capacity of a point-to-point DMC, which aligns with the lower bound established by Nascimento and Winter [Winter1]. Furthermore, when reduced to a specific erasure channel, they demonstrated a lower bound on the OT capacity. Thus, it remains uncertain what the exact OT capacity of a noisy DMC is in general. In practice, for general channels, a potential way is to first convert the channel into a Generalized Erasure Channel (GEC) via alphabet extension and erasure emulation, followed by the application of a general construction for GEC. As a starting point, we begin with a protocol by Ahlswede and Csiszár originated from the general two-party protocol of Section II:
@@
Two-Party OT Protocol [Rudolf1]: Consider the following two-party secure computation in the sense of OT: Alice has two strings $M_{0}$ and $M_{1}$ and aims to send them over the noisy point-to-point channel $\mathcal{W}:\mathcal{X}\rightarrow\mathcal{Y}$ to Bob. Bob has to choose one of them by inputting a bit of $Z\in\{0,1\}$ to the channel. Alice should be unaware of the unselected string, while Bob has only one string at the end of the protocol $(M_{Z})$. Suppose the main channel is an erasure channel assisted by a noiseless channel with unlimited capacity. The OT capacity in this setup is given by $\min(p,1-p)$, where $p$ is the erasure probability [Rudolf1, Imai]. Let $r<\min(p,1-p)$. The protocol by Ahlswede and Csiszár [Rudolf1], based on a technique originally introduced for a BSC in [Crepeau2, Sec. 6.4], proceeds as follows: Alice starts by transmitting a sequence $\mathbf{X}=X^{n}\sim\text{Bernoulli}(\frac{1}{2})$ of i.i.d. bits over the channel. Bob observes the channel output $\mathbf{Y}$. Let $E$ denote the set of indices at which $\mathbf{Y}$ is erased, and let $\overline{E}$ represent the set of indices where $\mathbf{Y}$ is not erased. If $|E|<nr$ or $|\overline{E}|<nr$, Bob aborts the protocol, as there are not enough erased or unerased bits to complete the protocol. From $\overline{E}$, Bob randomly selects a subset $S_{Z}$ of cardinality $nr$. From $E$, Bob randomly picks a subset $S_{\overline{Z}}$, also of size $nr$. Bob then shares the sets $S_{0}$ and $S_{1}$ with Alice via the public channel, where $S_{0}$ and $S_{1}$ are either $S_{Z}$ and $S_{\overline{Z}}$, respectively or vice versa. Alice cannot determine which of $S_{0}$ and $S_{1}$ corresponds to $E$ (erased positions) and which to $\overline{E}$ (non-erased positions) due to the independent nature of channel erasures. Using $S_{0}$ and $S_{1}$, Alice computes keys $\mathbf{X}|_{S_{0}}$ and $\mathbf{X}|_{S_{1}}$ and employs these keys to encrypt her strings, which she sends to Bob over the public channel as $M_{0}\oplus\mathbf{X}|_{S_{0}}$ and $M_{1}\oplus\mathbf{X}|_{S_{1}}$. Bob, who only knows the sequence corresponding to $\overline{E}$, can decrypt only one of these encrypted strings, depending on whether $S_{Z}=S_{0}$ or $S_{Z}=S_{1}$. This enables Bob to retrieve one of the two keys, $M_{Z}$, while he learns nothing about the other key, $M_{\overline{Z}}$. If $X$ is not uniformly distributed over $\{0,1\}$, the strings $\mathbf{X}|_{S_{j}},j\in\{0,1\}$ are not directly suitable as encryption keys. They need to be transformed into binary strings of length $k<nr$ with a distribution approximately uniform over $\{0,1\}^{k}$. It is well-known that for any $\delta>0$, when $n$ is large, there exists a mapping $\kappa:\{0,1\}^{nr}\rightarrow\{0,1\}^{k}$ with $k=n(H(X)-\delta)$ such that $k-H(\kappa(X^{n}))$ is exponentially small.
@@
In [chou], the author extends the above protocol to pairwise oblivious transfer over a noiseless binary adder channel involving two senders and one receiver, assuming they are non-colluding and honest-but-curious. Each sender has two strings, and Bob has to choose one string from each sender while the unselected strings are hidden from his view. In this system, the output is defined as the sum of the inputs, $Y=X_{1}+X_{2}$, commonly referred to as the Binary Erasure Multiple Access Channel (BE-MAC). This channel uniquely determines the inputs except when they differ, in which case one can not identify the inputs with certainty, effectively resulting in an erasure. Specifically, erasures occur in two out of four possible input scenarios. The OT capacity of this channel is shown to be $R_{1}+R_{2}\leq\max_{P_{X_{1}}P_{X_{2}}}H(X_{1},X_{2}|Y)=\frac{1}{2}$.
@@ IV System Model
A key concept when analyzing noisy channels is the idea of redundant symbols [Winter-BC]. We have the following definition for DM-MACs, presented in [Winter1] for the point-to-point channel.
@@
As proved in [Winter1] for point-to-point channels, a perfect DM-MAC (even after removing the redundancy) cannot be used for oblivious transfer, even against passive adversaries. This relates to the concept of noise and the emergence of noisy resources for cryptographic intents. We know that the noise produces uncertainty or entropy. It is the unpredictable aspect of a system that adds complexity. Conceptually, it can be seen as the manifestation of disorder or randomness in communication systems, emphasizing the non-perfect nature of the real world. Noise plays a crucial role in securing communication by hindering an eavesdropper’s ability to extract meaningful information from the transmitted message. In a noisy channel, the inherent noise limits the amount of information an unauthorized party can access, regardless of their computational power. This concept is central to Wyner’s wiretap channel model [Wyner], which demonstrates how noise can be leveraged to ensure that the legitimate receiver decodes the message accurately. In contrast, experiencing additional noise, the eavesdropper cannot gather sufficient information to reconstruct the message. As is clear, a perfect channel can be simulated by a noiseless channel where the input(s) can be obtained with certainty from the channel output(s). Obviously, such channels cannot be used for cryptographic intents with unconditional security (information-theoretic secrecy).
As is proved in [Imai, Rudolf1], the OT capacity of the point-to-point erasure channel (BEC) with erasure probability $\frac{1}{2}$ is equal $\frac{1}{2}$. Here, we want to investigate whether the OT is possible over a special BE-MAC. We introduce the channel as a correlation between the senders (Alice-1, Alice-2) and the receiver (Bob). Before that, we delve deeper into the unfairness in the channel/correlation model. Damgård et. al. [Damgard], introduced unfairness so that an unfair player could change the communication channel parameters within a certain range. In [Winter1], this concept is limited so that an unfair player who deviates from the channel statistics in $\delta n$ positions will be caught by the other party with probability $\geq 1-C_{1}\exp(-C_{2}\delta^{2}n)$, where $C_{1}$ and $C_{2}$ are two small positive numbers. There are two senders in our channel model. The concept of unfairness can also be extended to the MAC model. To control the fairness of other players in the last $n$ rounds, all players have access to a test unit. When both senders send a pair symbol to Bob over the channel, Bob will ask for the input symbols with probability $\frac{1}{2}$. He will tell both senders his output when he has received the senders’ response. If Alice-$i$ is unfair, she tells her input wrong; if Bob is unfair, he tells his output wrong. The test between players is to check out the samples after $n$ uses of the channel for joint typicality relative to $P_{X_{1}X_{2}Y}$.
@@
Our protocol to achieve $\text{SU-SBC}_{p,W}$ closely follows the approach outlined in [Winter1, Crepeau1]. It uses the source coding protocol $P_{X_{1}X_{2}Y}$ twice for two independent realizations. Let:
@@
Bob cannot behave unfairly beyond what is provided by the SUCO $P_{X_{1}X_{2}Y}$. However, the senders can introduce bias by attempting actions such as repeating unfair pairs $(x_{1},x_{2})(x_{1},x^{\prime}_{2}),(x_{1},x_{2})(x^{\prime}_{1},x_{2})$ or $(x_{1},x_{2})(x^{\prime}_{1},x^{\prime}_{2})$ for a good pair $i\in\mathcal{L}$. Suppose the senders persist with such strategies for $\delta n$ times (where $\delta>0$). In that case, Bob can detect their bias by approximating a typicality test, using Lemma 6 and an extended version of [Winter1, Lemma 6]. This aligns with the definition of $\text{SBC}_{p,W}$.
@@ V-A1 Information-Theoretic Formulation of OT
In the foundational works [Blundo, Nikov], the information-theoretic security definition for the receiver’s privacy requires that the sender’s final view, denoted by $U=(U^{\prime},R_{A},\mathbf{C})$, be statistically independent of the receiver’s choice bit $z\in\{0,1\}$. Here, $U^{\prime}$ represents the sender’s initial state (including inputs), $R_{A}$ denotes the sender’s private randomness, and $\mathbf{C}$ captures the complete public communication between the sender and receiver. However, as highlighted by Crépeau in [Crepeau], such a strong condition is generally unachievable in many practical settings. In particular, when there is a known dependency between the parties’ inputs, the sender’s view will unavoidably correlate with the receiver’s input. This is because the sender’s input $X$, which is inherently part of her view, may be statistically dependent on the receiver’s input $Z$. To address this, a more appropriate formulation requires that the sender’s view be conditionally independent of the receiver’s input, given the sender’s input. Formally, this is expressed via the Markov chain $(U^{\prime},R_{A},\mathbf{C})\rightarrow X\rightarrow Z$, which implies the conditional mutual information satisfies $I(U^{\prime},R_{A},\mathbf{C};Z\mid X)=0$. Nevertheless, Ahlswede and Csiszár [Rudolf1] adopt a slightly relaxed criterion, aligning with the interpretation in [Blundo, Nikov], where the privacy condition is written as $I(M_{0},M_{1},X^{n},R_{A},\mathbf{C};Z)\rightarrow 0$. This formulation is weaker, as it allows for some residual dependence under the assumption that it diminishes asymptotically.
@@
Formulating the security definition for the sender presents additional challenges beyond those encountered in defining receiver security [Blundo, Nikov]. Several issues commonly arise in this context:
@@
Notably, in [Brassard1, Darco], the model constrains a malicious receiver to alter their input in a purely deterministic manner. That is, the effective input $Z^{\prime}$—representing the bit ultimately learned by the receiver—is required to be a deterministic function of their original input $Z$. This stands in contrast to the ideal functionality of oblivious transfer, where a malicious receiver is permitted to choose $Z^{\prime}$ probabilistically, potentially based on arbitrary strategies. Such a deterministic restriction limits the adversary’s capabilities in the simulation, potentially weakening the security guarantee for the sender in realistic adversarial settings.
@@
In scenarios where a malicious receiver may alter their effective input from $z$ to $z^{\prime}$ (with $z,z^{\prime}\in\{0,1\}$), [Brassard2] allows $Z^{\prime}$ to depend on the sender’s input $X$. Such dependency is not permitted in the ideal model, where parties’ inputs and outputs are generated independently of each other through a trusted functionality. Furthermore, the security definition requires that the view of the malicious receiver—denoted by $(V^{\prime},R_{B},Y^{n},\mathbf{C})$—be conditionally independent of the sender’s input $X$, given the receiver’s original input $Z$ and their output $M_{Z^{{}^{\prime}}}$. This corresponds to the Markov chain: $X\rightarrow(Z,M_{Z^{{}^{\prime}}})\rightarrow(V^{\prime},R_{B},Y^{n},\mathbf{C}),$ which implies: $I(V^{\prime},R_{B},Y^{n},\mathbf{C};X\mid Z,M_{Z^{{}^{\prime}}})=0$.
@@
However, a more accurate formulation would account for the potential dependence of the receiver’s strategy on $Z^{\prime}$, leading to the stronger Markov condition: $X\rightarrow(Z,Z^{\prime},M_{Z^{{}^{\prime}}})\rightarrow(V^{\prime},R_{B},Y^{n},\mathbf{C})$. Ahlswede and Csiszár [Rudolf1] propose an alternative criterion under the honest-but-curious model—where Alice and Bob follow the protocol faithfully—that requires:
@@
In the context of OT, a malicious receiver may alter their effective input from $z$ to $z^{\prime}$ for $z,z^{\prime}\in\{0,1\}$. In [Brassard2], the variable $Z^{\prime}$—representing the receiver’s effective choice—may depend on the honest sender’s input $X$, which deviates from the ideal functionality where such a dependency is disallowed. Moreover, the receiver’s view, denoted by $(V^{\prime},R_{B},Y^{n},\mathbf{C})$, is required to be conditionally independent of the sender’s input $X$, given the receiver’s original input $Z$ and the corresponding output $M_{Z^{{}^{\prime}}}$. This condition corresponds to the Markov chain: $X\rightarrow(Z,M_{Z^{{}^{\prime}}})\rightarrow(V^{\prime},R_{B},Y^{n},\mathbf{C})$, or equivalently, $I(V^{\prime},R_{B},Y^{n},\mathbf{C};X\mid Z,M_{Z^{{}^{\prime}}})=0$. However, a more accurate and robust formulation accounts for the dependency of the adversary’s strategy on the full pair $(Z,Z^{\prime})$, implying the stronger Markov chain: $X\rightarrow(Z,Z^{\prime},M_{Z^{{}^{\prime}}})\rightarrow(V^{\prime},R_{B},Y^{n},\mathbf{C})$. Ahlswede and Csiszár [Rudolf1] propose a different formulation in the honest-but-curious setting, where both parties follow the protocol faithfully. In this model, Alice’s input is the message pair $(M_{0},M_{1})$, and Bob’s input is $Z$. Their security criterion requires that: $I(Z,R_{B},Y^{n},\mathbf{C};X)\triangleq I(Z,R_{B},Y^{n},\mathbf{C};M_{\overline{Z}})$, and they study its asymptotic behavior under the i.i.d. assumption. While this approach simplifies the analysis for semi-honest participants, it is too permissive under malicious adversaries and may inadvertently allow protocols with security vulnerabilities.